Question text

Chongo Co. had the following balances reported on their December 31st 20x9 financial statements:

Mortgage payable

$911987

Total Assets

$4049994

Long-term bank loan

$815132

Accounts payable

$543170

Other current liabilities

$217073


What is the debt-to-equity ratio for Chongo Co. using the information above? Debt is defined as long-term liabilities only. All of the company’s liabilities are listed above.

Select one:

a. 1.59

b. 1.11

c. 0.43

d. 0.61
Answer #1

Calculation of debt equity ratio:
Debt equity ratio= debt/equity
Debt= mortgage payable+ long term bank loan= 911987+815132= $1727119

Equity= Total assets-(mortgage payable+accounts payable+other current liabilities+long term bank loan)

Equity= 4049994-(911987+543170+217073+815132)= $1562632

Debt equity ratio= 1727119/1562632= 1.11
So correct answer is b) 1.11